OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) + 6

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

do you know the answer ??

OpenStudy (skullpatrol):

Yes.

OpenStudy (anonymous):

well could you please give it to me ?? its important

OpenStudy (anonymous):

?

OpenStudy (skullpatrol):

The answer is the number that when you replace x with it you will get a true statement using $$3(x + 1) = −2(x − 1) + 6$$

OpenStudy (anonymous):

ok what is it cause you are not helping me at all ??

OpenStudy (skullpatrol):

What grade math is this for?

OpenStudy (anonymous):

math 13

OpenStudy (anonymous):

please give me the answer i need it fast

mathslover (mathslover):

@msellier , Sorry but this site does NOT offer direct answers. I request you to let us know, what you tried and we will guide you accordingly. Please have a look to the OpenStudy code of conduct : http://openstudy.com/code-of-conduct

mathslover (mathslover):

@msellier , it a test question? If yes, then please review the OpenStudy NON CHEATING POLICY : http://openstudy.com/terms-and-conditions

OpenStudy (anonymous):

ok i get it sorry

OpenStudy (anonymous):

so could you help me explain me how to do this problem Solve for x: 3(x + 1) = −2(x − 1) + 6

mathslover (mathslover):

@skullpatrol , please continue! Thanks for patience.

mathslover (mathslover):

@msellier , first of all. Expand 3(x+1) , can you do that?

OpenStudy (anonymous):

3x + 3

mathslover (mathslover):

Good. Now expand -2(x-1)

OpenStudy (anonymous):

-2x + 2

OpenStudy (anonymous):

then i dont know what to do ?

mathslover (mathslover):

Now you have : 3x + 3 = -2x+ 2 + 6 Can you transpose -2x to LHS (Left Hand Side) ?

OpenStudy (anonymous):

um... then i dont have a clue of what i should do

mathslover (mathslover):

See, try to add (-2x) both sides. Can you do that?

OpenStudy (anonymous):

oh no i see 3x + 3 = -2x+ 2 + 6 3= -5x +2 + 6 3= -5x + 8 -5 divied by -5x which will give 1 is that write

OpenStudy (anonymous):

right*

mathslover (mathslover):

excellent.

mathslover (mathslover):

good work @msellier .

OpenStudy (anonymous):

ok thank you

mathslover (mathslover):

Did you like OpenStudy ?

OpenStudy (anonymous):

yes its good for helping i heard

mathslover (mathslover):

Msellier, I want to thank you first, you were a new user and I just warned you. You took it positively and see how it helped. If I had given you direct answer, then it will have been a demerit for you as you will have followed the direct answer method. Hope you understand and keep on going on this site.

OpenStudy (anonymous):

i have to go but i willl be back in an hour and 3O min to get more help for other problems

OpenStudy (anonymous):

i understood thank you